# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ F.4 Sea Temperature Prediction
F.5 Robotic Task Reconstruction
BridgeData V2 (walke2023bridgedata) is a large and diverse dataset of robotic manipulation behaviors, designed to advance research in scalable robot learning. In this experiment, our goal is to reconstruct the full video of a robot performing manipulation tasks while training the network only on images from the beginning and end of the sequence, interpreted as samples from the endpoint distributions $\rho_{a}$ and $\rho_{b}$. No intermediate marginals $\rho_{m}$ are used. Following the Sea Temperature Prediction experiment, we introduce a potential function $U$ that penalizes deviations from the learned data manifold. This manifold is learned from the initial and final frames of the videos, and the penalty encourages plausible intermediate frames consistent with these distributions. We employ a state-of-the-art VAE architecture, with parameters listed in Table 31, to model the image manifold. The potential function $U(x^{t})$ for samples $x^{t}\sim\rho^{t}$ is defined as the squared distance between a bridge sample $x^{t}$ and its VAE reconstruction $\tilde{x}^{t}=\text{VAE}(x^{t})$: $U(x^{t})=\lVert x^{t}-\tilde{x}^{t}\rVert^{2}$. Fig. 7 presents snapshots of the reconstructions produced by our CWG method compared to the baselines. In this experiment, the Hamiltonian function $H^{t_{k}}$ is held constant, as varying it yields no apparent benefit.
@@ F.5 Robotic Task Reconstruction
Figure 14: Reconstructions from CWG (top), GSBM (middle), and DSBM (bottom) in the Robot Task Reconstruction experiment. Red row shows the reference.
Table 31: Architecture of the ConvVAE used in the Robot Task Reconstruction experiment. All Conv2D and ConvTranspose2D layers use kernel size 44, stride 22, and padding 11, followed by ReLU activations (except the last layer, which uses Sigmoid).
@@ F.6 Unpaired Image Transfer
The Flickr-Faces-HQ (FFHQ) dataset is a high-resolution ($1024\times 1024$) collection of human faces that exhibits a remarkably wide range of visual variations (karras2019style). We closely follow the experimental protocol adopted by gushchin2024light, using their publicly available datasets from https://github.com/ngushchin/LightSB. We split the dataset into training (first $60K$ images) and testing (last $10K$ images) subsets. Each subset is further partitioned into age groups, specifically adults and children, corresponding to the source distribution $\rho_{a}$ and the target distribution $\rho_{b}$, respectively. The objective of this experiment is to generate a realistic child image from a given adult image, representing the same individual at a younger age. For each image, we employ the pre-trained ALAE encoder (pidhorskyi2020adversarial) to extract a $512$-dimensional latent vector. The SB solvers are trained directly on these latent representations. During inference, we encode a given image into the latent space, apply the learned bridge using the SB models, and then decode its outcome to obtain the transformed image. Since the probabilistic bridge is already defined in the latent space, there is no need for a potential function $U$ to measure the distance from the data manifold. Consequently, in the CWG framework we set $U=0$ and solve a constant-kinetic-energy bridge problem. In contrast, in the baseline GSBM approach we retain a potential energy term $U$ that penalizes excessive entropy and overly concentrated probability densities $\rho^{t}$, unlike DSBM, which does not include any potential term $U$.
@@
Qualitative examples of the images generated by our CWG method and the baseline models are presented in Figures 1517. A more compact, side-by-side qualitative comparison is provided in Figure 9, while quantitative results are reported in Table 7. The Feasibility metric reported in the table indicates whether the computed bridges respect the imposed constraints, that is, whether they match the target marginal distribution. It is computed by evaluating the FID score between the terminal distribution generated by the models, $\rho^{t_{K}}_{\theta}$, and the target distribution, $\rho_{b}$, as defined in equation (66b). The starting distribution $\rho_{a}$ corresponds to the source domain from which the input samples are drawn, and therefore it is naturally satisfied and does not require matching. The Optimality metric measures the geodesic length of the Schrödinger bridges computed by the models, in order to assess the transport cost that characterizes each solution. It is computed by sampling from the probabilistic path $z^{t_{k}}\sim\rho^{t_{k}}$ in the latent space. The pre-trained ALAE decoder is then used to map these latent vectors to images $x^{t_{k}}$, which are subsequently encoded into feature representations $x_{\text{fea}}^{t_{k}}$ using a pre-trained Inception-v3 network (the same used to compute the FID score). The curve length is then obtained from equation (66a).
Optimality $=d_{\text{feature}}=\sum_{k=1}^{K}d_{\mathcal{W}_{2}}(x_{\text{fea}}^{t_{k}},x_{\text{fea}}^{t_{k-1}});$ (66a)
@@
Hence, the length of the geodesic computed from the bridge in the latent space is evaluated according to the Wasserstein metric, pulled back from the feature space to the image space, and from the image space to the latent space. Although the baseline models achieve better Optimality scores (lower values), this improvement comes at the cost of failing to satisfy the marginal constraints, as reflected by their poorer Feasibility scores. Since constraint satisfaction is of critical importance, our CWG method exhibits superior overall behavior. This conclusion is further supported by Figure 18, where three facial age estimation networks, $h_{1}$ (prithivMLmods/facial-age-detection), $h_{2}$ (nateraw/vit-age-classifier), and $h_{3}$ (abhilash88/age-gender-prediction), were used to estimate the ages of the generated faces at each generation step. The different curves in the plot represent the predicted ages for the images shown in Figures 1517. While the CWG model exhibits a more decisive transformation across time steps, the baseline curves remain comparatively flat, often producing outputs that fail to convincingly resemble children. The age prediction models ($h_{1}$, $h_{2}$, $h_{3}$) provide stochastic metrics that quantify their predictive uncertainty, which we use to estimate their variance $\sigma^{2}$ under the assumption of a uniform error distribution. The average predicted age and the corresponding variance for the plots represented in Figure 18 are computed as,
@@
Table 33: FID scores ($\downarrow$) measuring the distance to the initial adult distribution $\rho_{a}$ and the target child distribution $\rho_{b}$ for samples generated at different time steps using the CWG method in the Unpaired Image Transfer.
@@
Figure 15: Adult $\rightarrow$ Child image generation in the Unpaired Image Transfer using the CWG method.
Figure 16: Adult $\rightarrow$ Child image generation Unpaired Image Transfer using the GSBM method.
Figure 17: Adult $\rightarrow$ Child image generation Unpaired Image Transfer using the DSBM method.
@@
Figure 18: Age prediction for the samples generated in the Unpaired Image Transfer experiment.

### v3

@@ G.4 Sea Temperature Prediction
G.5 Robotic Task Reconstruction
BridgeData V2 (Walke et al., 2023) is a large and diverse dataset of robotic manipulation behaviors, designed to advance research in scalable robot learning. In this experiment, our goal is to reconstruct the full video of a robot performing manipulation tasks while training the network only on images from the beginning and end of the sequence, interpreted as samples from the endpoint distributions $\rho_{a}$ and $\rho_{b}$. No intermediate marginals $\rho_{m}$ are used. Following the Sea Temperature Prediction experiment, we introduce a potential function $U$ that penalizes deviations from the learned data manifold. This manifold is learned from the initial and final frames of the videos, and the penalty encourages plausible intermediate frames consistent with these distributions. We employ a state-of-the-art VAE architecture, with parameters listed in Table 33, to model the image manifold. The potential function $U(x^{t})$ for samples $x^{t}\sim\rho^{t}$ is defined as the squared distance between a bridge sample $x^{t}$ and its VAE reconstruction $\tilde{x}^{t}=\text{VAE}(x^{t})$: $U(x^{t})=\lVert x^{t}-\tilde{x}^{t}\rVert^{2}$. Fig. 7 presents snapshots of the reconstructions produced by our CWG method compared to the baselines. In this experiment, the Hamiltonian function $H^{t_{k}}$ is held constant, as varying it yields no apparent benefit.
@@ G.5 Robotic Task Reconstruction
Figure 14: Reconstructions from CWG (top), GSBM (second), DSBM (third), and SB-Flow (bottom) in the Robot Task Reconstruction experiment. Red row shows the reference.
Metric CWG (ours) GSBM DSBM SB-Flow
FID $\bm{{18.83}_{\pm 0.66}}$ ${40.23}_{\pm 1.95}$ ${149.78}_{\pm 0.81}$ ${73.46}_{\pm 0.52}$
tt (s) $\bm{{1090}_{\pm 40}}$ ${91100}_{\pm 8000}$ ${27400}_{\pm 2500}$ ${4900}_{\pm 300}$
Table 33: Architecture of the ConvVAE used in the Robot Task Reconstruction experiment. All Conv2D and ConvTranspose2D layers use kernel size $4$, stride 22, and padding 11, followed by ReLU activations (except the last layer, which uses Sigmoid).
@@ G.6 FFHQ Transfer
The Flickr-Faces-HQ (FFHQ) dataset is a high-resolution ($1024\times 1024$) collection of human faces that exhibits a remarkably wide range of visual variations (Karras et al., 2019). We closely follow the experimental protocol adopted by Gushchin et al. (2024), using their publicly available datasets from https://github.com/ngushchin/LightSB. We split the dataset into training (first $60K$ images) and testing (last $10K$ images) subsets. Each subset is further partitioned into age groups, specifically adults and children, corresponding to the source distribution $\rho_{a}$ and the target distribution $\rho_{b}$, respectively. The objective of this experiment is to generate a realistic child image from a given adult image, representing the same individual at a younger age. For each image, we employ the pre-trained ALAE encoder (Pidhorskyi et al., 2020) to extract a $512$-dimensional latent vector. The SB solvers are trained directly on these latent representations. During inference, we encode a given image into the latent space, apply the learned bridge using the SB models, and then decode its outcome to obtain the transformed image. Because the probabilistic bridge is already defined within the ALAE latent space, which captures the FFHQ data manifold, the CWG method does not rely on the potential function $U$ to quantify deviations from image manifold. In contrast, the baseline GSBM approach retains a potential energy term $U$ that penalizes excessive entropy and overly concentrated probability densities $\rho^{t}$, distinguishing it from DSBM, which does not include any potential term.
@@
Qualitative examples of the images generated by our CWG method and the baseline models are presented in Figures 1517. A more compact, side-by-side qualitative comparison is provided in Figure 9, while quantitative results are reported in Table 7. The Feasibility metric indicates whether the computed bridges respect the imposed constraints, that is, whether they match the target marginal distribution. It is computed by evaluating the FID score between the terminal distribution generated by the models, $\rho^{t_{K}}_{\theta}$, and the target distribution, $\rho_{b}$, as defined in equation (72b). The starting distribution $\rho_{a}$ corresponds to the source domain from which the input samples are drawn, and therefore it is naturally satisfied and does not require matching. The Optimality metric measures the geodesic length of the Schrödinger bridges computed by the models, in order to assess the transport cost that characterizes each solution. It is computed by sampling from the probabilistic path $z^{t_{k}}\sim\rho^{t_{k}}$ in the latent space. The pre-trained ALAE decoder is then used to map these latent vectors to images $x^{t_{k}}$, which are subsequently encoded into feature representations $x_{\text{fea}}^{t_{k}}$ using a pre-trained Inception-v3 network (the same used to compute the FID score). The curve length is then obtained from equation (72a).
Optimality $=\sum_{k=1}^{K}d_{\mathcal{W}_{2}}(x_{\text{fea}}^{t_{k}},x_{\text{fea}}^{t_{k-1}});$ (72a)
@@
Hence, the length of the geodesic computed from the bridge in the latent space is evaluated according to the Wasserstein metric, pulled back from the feature space to the image space, and from the image space to the latent space. Although the baseline models achieve better Optimality scores (lower values), this improvement comes at the cost of failing to satisfy the marginal constraints, as reflected by their poorer Feasibility scores. Since constraint satisfaction is of critical importance, our CWG method exhibits superior overall behavior. This conclusion is further supported by Figure 19, where three facial age estimation networks, $h_{1}$ (prithivMLmods/facial-age-detection), $h_{2}$ (nateraw/vit-age-classifier), and $h_{3}$ (abhilash88/age-gender-prediction), were used to estimate the ages of the generated faces at each generation step. The different curves in the plot represent the predicted ages for the images shown in Figures 1517. While the CWG model exhibits a more decisive transformation across time steps, the baseline curves remain comparatively flat, often producing outputs that fail to convincingly resemble children images. The age prediction models ($h_{1}$, $h_{2}$, $h_{3}$) provide stochastic metrics that quantify their predictive uncertainty, which we use to estimate their variance $\sigma^{2}$ under the assumption of a uniform error distribution. The average predicted age and the corresponding variance for the plots represented in Figure 19 are computed as,
@@
Table 35: FID scores ($\downarrow$) measuring the distance to the initial adult distribution $\rho_{a}$ and the target child distribution $\rho_{b}$ for samples generated at different time steps using the CWG method in the FFHQ transfer.
@@
Figure 15: Adult $\rightarrow$ Child image generation in the FFHQ transfer using the CWG method.
Figure 16: Adult $\rightarrow$ Child image generation FFHQ transfer using the GSBM method.
Figure 17: Adult $\rightarrow$ Child image generation FFHQ transfer using the DSBM method.
Figure 18: Adult $\rightarrow$ Child image generation FFHQ transfer using the SB-Flow method.
@@
Figure 19: Age prediction for the samples generated in the FFHQ transfer experiment.
G.7 MNIST-to-EMNIST
The MNIST and EMNIST datasets are low-resolution ($28\times 28$) handwritten character datasets that serve as long-standing benchmarks for generative modeling, representation learning, and probabilistic transport (Cohen et al., 2017). In this work, we adopt MNIST as the source distribution $\rho_{a}$ and EMNIST (letters) as the target $\rho_{b}$ to study unpaired image-to-image translation. We restrict the analysis to ten digits and ten letters. The two datasets are visually related but statistically distinct, making the MNIST-to-EMNIST transport problem sufficiently non-trivial while still permitting clear qualitative and quantitative evaluation. Our goal is to produce terminal samples that match the EMNIST distribution while ensuring that the Schrödinger Bridge generates meaningful intermediate states that remain on the data manifold rather than collapsing into noise or artifacts.
For this purpose, we introduce a potential function $U$ that penalizes deviations from the learned data manifold. Building on the approach of Song and Itti (2025), where generative models are evaluated by measuring the distance between their outputs and a geometric manifold of real images learned by a VAE, we adopt a similar strategy. Specifically, we use a state-of-the-art VAE architecture, with parameters listed in Table 37, to learn a manifold composed of images from both marginal distributions $\rho_{a}$ and $\rho_{b}$. The potential function $U(x^{t})$, for samples $x^{t}\sim{\rho}^{t}$, is then defined as the squared distance between a bridge sample $x^{t}$ and its VAE-projected reconstruction $\tilde{x}^{t}=\text{VAE}(x^{t})$: $U(x^{t})=\lVert x^{t}-\tilde{x}^{t}\rVert^{2}$. No intermediate samples $\rho_{m}$ are used in this experiment.
The results are shown visually in Figure 20 and quantitatively in Table 38. As observed in the other experiments, DSBM fails to regularize the stochastic path between the two distributions, and SB-Flow searches for an interpolation that first removes features of the source digit not present in the target letter, rather than producing intermediate characters that remain close to the data manifold at every time step. GSBM produces only a deterministic trajectory along the data manifold, which it then uses as the mean of a transient Gaussian distribution. However, because the stochastic component is not properly constrained, these Gaussian samples can drift off the manifold, leading to unrealistic intermediate states. This drift appears as increased noise in the midpoint samples and results in higher FID scores. In contrast, the CWG method produces intermediate characters that remain closest to the source and target distributions at every discretized time step. Notably, our method also demonstrates strong computational efficiency, achieving training times significantly lower than those of the baselines.
To evaluate the guided version of the CWG method, we propose a Gaussian deblurring test, a well-established inverse problem commonly used in image restoration (Feng et al., 2025; Wang et al., 2025). In this setup, a reference image from the EMNIST dataset is processed with a Gaussian kernel and provided as the target for restoration. In this test, the guided loss function, $\|f({x}^{t_{K}})-y\|$, uses the blurry image as the target $y$, and the feature function $f$ is defined as,
$f({x}^{t_{K}})=\mathbf{K}_{\text{gaussian}}(\sigma,r)\odot{x}^{t_{K}},$ (74)
where $\mathbf{K}_{\text{gaussian}}(\sigma,r)$ is a Gaussian kernel operator of variance $\sigma$ and size $r$, applied to the sample ${x}^{t_{K}}$ generated at the last time step of the bridge. The specific kernel parameters are detailed in Table 36. This loss measures the similarity between the blurred version of the generated image and the blurred reference, ensuring compatibility between the two. The restoration results are shown visually in Figure 21 and quantitatively in Table 39.
Table 36: Parameters for the feature function $f$ (74), used for guided generation in the Mnist-to-Emnist experiment.
Parameter Value
Kernel variance $\sigma$ 33
Kernel size $r$ 99
Table 37: Architecture of the ConvVAE used in the MNIST-to-EMNIST experiment. All Conv2D and ConvTranspose2D layers use ReLU activations unless otherwise specified.
Stage Layer (channels) Output size
Input Single-channel image $1\times 28\times 28$
[…]
Flatten $128\cdot 7\cdot 7=6272$
Latent space Linear $\to\mu$ 22
Linear $\to\log\sigma^{2}$ 22
Decoder Linear $\to$ reshape $128\times 7\times 7$
[…]
Table 38: FID scores computed w.r.t. the EMNIST distribution ($\downarrow$) and training time (tt) in the MNIST-to-EMNIST experiment.
Metric CWG GSBM DSBM SB-Flow
[…]
tt (s) $\bm{570_{\pm 30}}$ $30100_{\pm 1000}$ $8750_{\pm 300}$ $2200_{\pm 60}$
(a) CWG (ours)
(b) GSBM
(c) DSBM
(d) SB-Flow
Figure 20: Snapshots from the bridges computed in the MNIST-to-EMNIST experiment.
Figure 21: Gaussian deblurring test for the Mnist-to-Emnist experiment. Results for the guided CWG method.
Metric Score
FID $12.72_{\pm 0.36}$}
LPIPS $0.28_{\pm 0.02}$}
PSNR $12.48_{\pm 1.11}$}
SSIM $0.37_{\pm 0.04}$}
tt (s) $320_{\pm 20}$}
Table 39: Quantitative metrics for the Gaussian deblurring experiment: FID score ($\downarrow$) between the clean and reconstructed samples, LPIPS ($\downarrow$), PSNR ($\uparrow$), SSIM $\uparrow$), training time (tt) ($\downarrow$).